Verify that f(z) is analytic and evaluatesz)dz : 10) je (11) ſz cos zdz
Answer #1

zezed ze d2 Let f(2)= zeze 2(22)+ e ci Them f(2)=ze - Plz) = ez(2271) - p2) #100 fa erosy zee >> 18 differentable at every se[zc8zdz tet f(2)=2682 Them f(2)=~25M2+QB2 s pcz)+00 rzec p is differentiate at every Zee R is cancelyter on a Zordz 2p78) -